# Dual chiral density wave induced oscillating Casimir effect

### v1

@@ IV Conclusion and outlook
  • Solving gap equation in finite volume—In this work, we input the values of order parameters obtained from the gap equation in infinite volume, in order to understand the typical behavior of the Casimir effect for typical dispersion relations. As a more precise analysis, one can calculate the order parameters by minimizing the thermodynamic or effective potential in finite volume, and then one can understand the self-consistent relationship between the volume-dependent order parameters and the Casimir effect. This type of study would help elucidate the phase diagram of interacting fermions in finite volume (e.g., see Refs. [100, 101, 102, 103, 104, 105, 106, 107]) from the viewpoint of the Casimir effect.
@@
[…]
@@ ACKNOWLEDGMENTS
Appendix A Results with MIT bag boundary
@@
The Lifshitz formula for the Casimir energy from the Dirac sea with the “MIT bag” boundaries is given by
$E_{\rm Cas}^{\rm``MIT"}$ $=-2N_{f}N_{c}\sum_{s=\pm}\int_{0}^{\infty}\frac{d\xi}{2\pi}\int\frac{d^{2}k_{\perp}}{(2\pi)^{2}}\ln\big{[}1+e^{-2L_{z}k_{z}^{[s]}}\big{]},$
$k^{[\pm]}_{z}$ $=\sqrt{k_{\perp}^{2}+M^{2}+\xi^{2}-\frac{q^{2}}{4}\mp iq\sqrt{k_{\perp}^{2}+\xi^{2}}},$ (20)
where the first integral variable is the imaginary part $\xi$ of imaginary frequency $\omega\equiv i\xi$, $k_{\perp}^{2}\equiv k_{x}^{2}+k_{y}^{2}$, and $d^{2}k_{\perp}\equiv dk_{x}dk_{y}$. The overall factor of $2N_{f}N_{c}$ means the degrees of freedom for particle-antiparticle, flavors, and colors, and the spin degrees of freedom are labeled as $s=\pm$. The overall minus sign and the factor of $+e^{-2L_{z}k_{z}^{[s]}}$ is a property of the “MIT bag” boundary on fermion fields. In the lattice rugularization, we have to replace as $\sum^{\rm BZ}_{n}\to\frac{1}{2}\sum^{\rm BZ}_{n}$ in Eq. (14), and the sum is over $n=0,1,...,2N_{z}-1$.
Figure 11: Casimir coefficients in the low-density region with the “MIT bag” boundary conditions. We denote $a\Lambda$ as $a$ in the legends.
Figure 11 shows the results of the Casimir coefficients at low-, intermediate-, and high-density regions. In the case of the “MIT bag” boundary, the number of discretized levels is twice larger than that with the PBC. As a result, the period of oscillation is half of that with the PBC: analogous to Eq. (19),
@@
Furthermore, from these results, we can also discuss the case with the antiperiodic boundary conditions (APBCs). The discrete momentum, $k_{z}\to(2n+1)\pi/L_{z}$, for the APBCs is shifted by a half period from $k_{z}\to 2n\pi/L_{z}$ for the PBCs, and hence the sign of the Casimir energy is flipped and also its absolute value changes slightly. The corresponding Lifshitz formula is written as
$E_{\rm Cas}^{\rm APBC}$ $=-4N_{f}N_{c}\sum_{s=\pm}\int_{0}^{\infty}\frac{d\xi}{2\pi}\int\frac{d^{2}k_{\perp}}{(2\pi)^{2}}\ln\big{[}1+e^{-L_{z}k_{z}^{[s]}}\big{]}.$ (22)
In the $3+1$ dimensions, the Casimir energy with $k_{z}\to(2n+1)\pi/L_{z}$ is $16$ times larger than that with $k_{z}\to(n+1/2)\pi/L_{z}$. As a result, by multiplying the vertical axis obtained in Fig. 11 by $16$ and the horizontal axis by $2$, we can obtain the exact results for the APBCs.

### v2

@@ IV Conclusion and outlook
  1. (i)
    Solving gap equation in finite volume—In this work, we input the values of order parameters obtained from the gap equation in infinite volume, in order to understand the typical behavior of the Casimir effect for typical dispersion relations. As a more precise analysis, one can calculate the order parameters by minimizing the thermodynamic or effective potential in finite volume, and then one can understand the self-consistent relationship between the volume-dependent order parameters and the Casimir effect. This type of study would help elucidate the phase diagram of interacting fermions in finite volume (e.g., see Refs. [100, 101, 102, 103, 104, 105, 106, 107]) from the viewpoint of the Casimir effect. [^7] As shown in Ref. [103], the phase diagram in the NJL model at finite $(\mu,L_{z})$ is complicated, where the phase boundary between low-density and high-density phases oscillates as a function of $L_{z}$. Therefore, for a more realistic prediction of the $L_{z}$ dependence of the Casimir energy, such $L_{z}$-dependent phase transitions should be considered. In addition, a question is whether or not the DCDW phase survives on the phase diagram. As a preliminary result, we solved a gap equation at finite $(\mu,L_{z})$ and confirmed that in fact the DCDW phase appears in a parameter region [108]. In this region, our prediction of the Casimir effect is qualitatively reliable.
  2. (ii)
@@
  7. (vii)
    Lower-dimensional models of QCD—The WP-induced oscillating Casimir effect can occur also in lower spatial dimensions (i.e., $1+1$ or $2+1$ dimensions) [31]. Therefore, investigation of the Casimir effect in possible inhomogeneous phases of low-dimensional QCD-like models, such as the Gross-Neveu model [130, 131, 132], the chiral Gross-Neveu (NJL2) model [133], and the ’t Hooft model (QCD2[133, 134, 135], will be also important.
@@ ACKNOWLEDGMENTS
Appendix A Lifshitz formula
In this Appendix, we derive the Lifshiz formula (11) for the dispersion relations in the DCDW phase. To calculate Casimir energy at zero temperature and zero chemical potential, we need to obtain the infinite sum of eigenvalues of quarks,
[…]
where the momentum in the $z$ direction is discretized as $k_{z}=2n\pi/L_{z}$ and $k_{z}=(2n+1)\pi/L_{z}$ under the periodic boundary conditions (PBCs) and the antiperiodic boundary conditions (APBCs), respectively. The prime in the sum for the PBC means that the factor $1/2$ is multiplied only for $n=0$: $\sum_{n}^{\prime}\omega_{n}=\sum_{n}\omega_{n}-\omega_{n=0}/2$.
The discretized $k_{z}$ is given as the zero point of the function
$\Delta_{\pm}^{\mathrm{PBC}}(\omega)=1-e^{-ik^{[\pm]}_{z}L_{z}},$ (21)
$\Delta_{\pm}^{\mathrm{APBC}}(\omega)=1+e^{-ik^{[\pm]}_{z}L_{z}},$ (22)
$ik^{[\pm]}_{z}\equiv\tilde{k}^{[\pm]}_{z}=\sqrt{k^{2}_{\perp}+M^{2}-\omega^{2}-\frac{q^{2}}{4}\mp iq\sqrt{k^{2}_{\perp}-\omega^{2}}},$
where the imaginary momentum $ik^{[\pm]}_{z}$ is obtained by solving the DCDW dispersion relation $\omega_{s}=\sqrt{k_{\perp}^{2}+(\sqrt{k^{2}_{z}+M^{2}}+sq/2)^{2}}$ with $s=\pm 1$. Since the function $\Delta_{s}(\omega)$ is a meromorphic function in a closed path $C$ on the complex $\omega$ plane with the mentioned zero point but no pole, by using the generalized argument principle [136, 137, 4], the infinite sum (20) can be calculated from the following contour integral:
$\sum_{s=\pm}\sum^{\infty}_{n=-\infty}\frac{\omega_{s,n}}{2}=\sum_{s=\pm}{\sum^{\infty}_{n=0}}^{\prime}\omega_{s,n}=\sum_{s=\pm}\frac{1}{2\pi i}\oint_{C}\omega d\ln\Delta_{s}(\omega)$
$=\sum_{s=\pm}\frac{1}{2\pi i}\Big{(}\int^{-i\infty}_{i\infty}\omega d\ln\Delta_{s}(\omega)+\int_{C_{+}}\omega d\ln\Delta_{s}(\omega)\Big{)},$ (23)
where $C_{+}$ is the counterclockwise integral on a semicircle with an infinite radius centered at the origin, in the right half of the complex $\omega$ plane. More precisely, when at $M=q=k_{\perp}=0$ under the PBCs, the zero mode exists at the origin, and then the contour needs to avoid the origin by considering an infinitesimal semicircular path centered at the origin. By introducing the imaginary frequency $\omega\equiv i\xi$ with the imaginary part $\xi$, we obtain
$\sum_{s=\pm}\sum^{\infty}_{n=-\infty}\frac{\omega_{s,n}}{2}$
$=\sum_{s=\pm}\frac{1}{2\pi}\Big{(}\int^{-\infty}_{\infty}\xi d\ln\Delta_{s}(i\xi)+\int_{C_{+}}\xi d\ln\Delta_{s}(i\xi)\Big{)}.$ (24)
The second term of Eq. (24) on a circle with an infinite radius vanishes in the limit of $\xi\to\infty$:
$\xi d\ln\Delta_{s}(i\xi)=\frac{L_{z}\xi^{2}e^{-\xi L_{z}}}{(1-e^{-\xi L_{z}})\xi}\xrightarrow[\xi\rightarrow\infty]{}0.$ (25)
Using integration by parts, the first term of Eq. (24) is reduced to
[…]
In the last form, we used $\xi\ln(1\pm e^{-\xi})\to 0$ in the limit of $\xi\to\infty$ and the fact that the integrand of the second term is an even function of $\xi$.
By taking into account the $k_{\perp}$ integral and multiplying the factors $2N_{f}N_{c}$ for the particle-antiparticle, flavors, and colors, we obtain the Lifshiz formula (11) for the PBC. Similarly, the Lifshitz formula for the APBC is
$E_{\rm Cas}^{\rm APBC}$ $=-4N_{f}N_{c}\sum_{s=\pm}\int_{0}^{\infty}\frac{d\xi}{2\pi}\int\frac{dk_{x}dk_{y}}{(2\pi)^{2}}\ln\big{[}1+e^{-L_{z}\tilde{k}_{z}^{[s]}}\big{]},$
$\tilde{k}^{[\pm]}_{z}$ $=\sqrt{k_{\perp}^{2}+M^{2}+\xi^{2}-\frac{q^{2}}{4}\mp iq\sqrt{k_{\perp}^{2}+\xi^{2}}}.$ (27)
The overall $-2$ and the factor of $+e^{-L_{z}\tilde{k}_{z}^{[s]}}$ is a property of the APBC on fermion fields. When we substitute $M=q=0$ and $q=0$, Eq. (27) returns to the known formulas for the massless and massive quarks, $E_{\rm Cas}^{\rm APBC}=-N_{f}N_{c}\times 7\pi^{2}/180L_{z}^{3}$ and $E_{\rm Cas}^{\rm APBC}=N_{f}N_{c}\times(-2M^{2}/\pi^{2}L_{z})\sum_{l=1}^{\infty}(-1)^{l+1}K_{2}(lML_{z})/l^{2}$, respectively. The discrete momentum, $k_{z}\to(2n+1)\pi/L_{z}$, for the APBCs is shifted by a half period from $k_{z}\to 2n\pi/L_{z}$ for the PBCs, and hence the sign of the Casimir energy is flipped and also its absolute value changes slightly.
Appendix B Results with MIT bag boundary
@@
Figure 11: Casimir coefficients in the zero-, intermediate-, and high-density regions with the “MIT bag” boundary conditions. We denote $a\Lambda$ as $a$ in the legends.
In this case, the infinite sum of eigenvalues is
$\sum_{s=\pm}{\sum^{\infty}_{n=0}}\frac{\omega_{s,n}^{\rm``MIT"}}{2},$ (28)
and, the discretized $k_{z}$ is given as the zero point of the function
$\Delta^{\rm``MIT"}_{\pm}(\omega)=1+e^{-2i\tilde{k}^{[\pm]}_{z}L_{z}}.$ (29)
Using the derivation in Appendix A, we obtain the Lifshitz formula with the “MIT bag” boundaries:
$E_{\rm Cas}^{\rm``MIT"}$ $=-2N_{f}N_{c}\sum_{s=\pm}\int_{0}^{\infty}\frac{d\xi}{2\pi}\int\frac{d^{2}k_{\perp}}{(2\pi)^{2}}\ln\big{[}1+e^{-2L_{z}\tilde{k}_{z}^{[s]}}\big{]},$
$\tilde{k}^{[\pm]}_{z}$ $=\sqrt{k_{\perp}^{2}+M^{2}+\xi^{2}-\frac{q^{2}}{4}\mp iq\sqrt{k_{\perp}^{2}+\xi^{2}}},$ (30)
where $k_{\perp}^{2}\equiv k_{x}^{2}+k_{y}^{2}$, and $d^{2}k_{\perp}\equiv dk_{x}dk_{y}$. The overall minus sign and the factor of $+e^{-2L_{z}\tilde{k}_{z}^{[s]}}$ is a property of the “MIT bag” boundary on fermion fields. In the lattice regularization, we have to replace as $\sum^{\rm BZ}_{n}\to\frac{1}{2}\sum^{\rm BZ}_{n}$ in Eq. (14), and the sum is over $n=0,1,...,2N_{z}-1$.
Figure 11 shows the results of the Casimir coefficients at zero-, intermediate-, and high-density regions. In the case of the “MIT bag” boundary, the number of discretized levels is twice larger than that with the PBC. As a result, the period of oscillation is half of that with the PBC: analogous to Eq. (19),
@@
Furthermore, from these results, we can also discuss the relationship between the cases with the “MIT bag” boundary and the APBC. The discrete momenta are $k_{z}\to(n+1/2)\pi/L_{z}$ and $k_{z}\to(2n+1)\pi/L_{z}$, respectively. For this reason, Eq. (30) is equal to Eq. (27) multiplying by $1/2$ and replacing as $L_{z}\to 2L_{z}$. As a result, the Casimir energy with “MIT bag” boundaries is exactly $16$ times smaller than that with the APBCs (the factor of $2^{3}$ is due to $3+1$ dimensions). Therefore, using the results in Fig. 11, by multiplying the vertical axis by $16$ and the horizontal axis by $2$, we can obtain the exact results for the APBCs.